OpenStudy (anonymous):

Find the absolute minimum and absolute maximum values of f on the interval below. f(x) = x^4 - 8x^2 + 8 with point[-3, 4]

OpenStudy (anonymous):

\[f(x) = x^4 - 8x^2 + 8\] \[f'(x)=4x^3-16x\] critical points at \[4x^3-16x=0\] \[4x(x^2-4)=0\] \[4x)x+2)(x-2)=0\] \[x=-2,0,2\]

OpenStudy (anonymous):

your last job is finding \[f(-3),f(-2),f(0),f(2),f(4)\] biggest is max, smallest is min